OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of two-thirds? 2x + 3y = 3 2x – 3y = 3 3x + 2y = 3 3x – 2y = 3

OpenStudy (anonymous):

y-3=2/3(x- 6) y= 2/3x -4 + 3 y= 2/3x -1 } x 3 3y = 2x -3 3y - 2x = -3 } x -1 2x - 3y =3

OpenStudy (anonymous):

Thank you :)